Given the population: 1, 3, 4, 6, and 8; and suppose samples of size 3 are drawn from this population:

1. What is the mean and standard deviation of the population?

2. How many different samples of size n=3 can be drawn from the population? List them with their corresponding means.

3. Construct the sampling distribution of the sample means.

4. What is the mean of the sampling distribution of the sample means? Compare this to the mean of the population.

5. What is the standard deviation of the sampling distribution of the sample means? Compare this to the standard deviation of the population.



1
Expert's answer
2021-03-25T01:58:49-0400

1. Mean


"\mu=\dfrac{1+3+4+6+8}{5}=4.4"

Variance

"\sigma^2=\dfrac{1}{5}\big((1-4.4)^2+(3-4.4)^2+(4-4.4)^2"

"(6-4.4)^2+(8-4.4)^2\big)=5.84"


Standard deviation


"\sigma=\sqrt{\sigma^2}=\sqrt{5.84}\approx2.4166"



2. We have population values "1,3,4,6,8," population size "N=5" and sample size "n=2." Thus, the number of possible samples which can be drawn without replacement is


"\dbinom{N}{n}=\dbinom{5}{2}=10""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 1,3,4 & 8/3 \\
 \hdashline
 2 & 1,3,6 & 10/3 \\
 \hdashline
 3 & 1,3,8 & 4 \\
 \hdashline
 4 & 1,4,6 & 11/3 \\
 \hdashline
 5 & 1,4,8 & 13/3 \\
 \hdashline
 6 & 1,6,8 & 5 \\
 \hdashline
 7 & 3,4,6 & 13/3 \\
 \hdashline
 8 & 3,4,8 & 5 \\
 \hdashline
 9 & 3,6,8 & 17/3 \\
 \hdashline
 10 & 4,6,8 & 6 \\ \hline
\end{array}"

3. The sampling distribution of the sample means.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 8/3 & 1& 1/10 & 8/30 & 64/90 \\
 \hdashline
 10/3 & 1& 1/10 & 10/30 & 100/90 \\
 \hdashline
 11/3 & 1& 1/10 & 11/30 & 121/90 \\
 \hdashline
 4 & 1& 1/10 & 12/30 & 144/90 \\
 \hdashline
 13/3 & 2& 2/10 & 26/30 & 338/90 \\
 \hdashline
 5 & 2& 2/10 & 30/30 & 450/90 \\
 \hdashline
 17/3 & 1& 1/10 & 17/30 & 289/90 \\
 \hdashline
 6 & 1& 1/10 & 18/30 & 324/90 \\
 \hdashline
 Total & 10 & 1 & 132/30 & 1830/90 \\ \hline
\end{array}"

4.


"E(\bar{X})=\sum\bar{X}f(\bar{X})=\dfrac{132}{30}=4.4"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.


"E(\bar{X})=4.4=\mu"

5.

"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"

"=\dfrac{1830}{90}-(\dfrac{132}{30})^2=\dfrac{876}{900}=\dfrac{73}{75}"

"\sqrt{Var(\bar{X})}=\sqrt{\dfrac{73}{75}}\approx0.9866"

Verification:

"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{5.84}{3}(\dfrac{5-3}{5-1})"

"=\dfrac{5.84}{6}=\dfrac{73}{75}, True"